Use a quadratic equation to find two real numbers with a sum of 2 and a product of -24.
Answer by Alan3354(69443)

Use a quadratic equation to find two real numbers with a sum of 2 and a product of -24.
---------------------
x + y = 2
x*y = -24
-------
x + y = 2 ---> y = 2-x
Sub for y in the 2nd eqn
==========
x*(2-x) = -24
x*(x-2) = 24
x^2 - 2x - 24 = 0
======
Here, find 2 factors of -24 that differ by 2
---> -6 & 4 : Obviously the "answer" -- no quadratic is needed.
----
(x-6)*(x+4) = 0
---> 6 and -4